Mikaela1509:

plz help me what is 60 times 50 divided by 3 and rounded?

snowflake0531:

60*50/3 What is 60*50

Mikaela1509:

um 300,000?.............

Mikaela1509:

maybe

snowflake0531:

What is 50*60 What is 5 times 6?

Mikaela1509:

30

snowflake0531:

So, What is 30 * 10*10

Mikaela1509:

3000?

snowflake0531:

Yes So what is 3000/3

Mikaela1509:

1000?

snowflake0531:

Yes~

Mikaela1509:

wait really?

snowflake0531:

Yes Another easy way to do 60*50/3 Is First, Divide 3 from 60, which is 20, then multiply it by 50

Mikaela1509:

oh ok thank you so much

MidNiGhtSkY:

and since it isn't a decimal you don't need to round

snowflake0531:

Yw~!

jhonyy9:

or an easier way to solve it \[\frac{ 60*50 }{ 3 } = \frac{ 3*20*50 }{ 3 } = 20*50 = ?\] 3 from numerator and 3 from denominator will simplified